OpenStudy (crisulcampo):

according to the purpose of deciding if lots of computers that sends a factory are accepted, a procedure is done that consist on choosing 10 equips radomly from each lot and determinate the number of obsolete ones. Any lot is always refused if it's found 3 or more obsolete equips among 10 chosen lots. Suposse that number of items in every lot is big and each lot contains 5% obsolete equips. ¿what is the probability of accepting a lot? @agent0smith @mathstudent55 @ganeshie8 Please guys, help with this one (five points and your medal)

OpenStudy (crisulcampo):

I don't know how to handle this one

OpenStudy (crisulcampo):

No ideas in my mind now

OpenStudy (crisulcampo):

@ganeshie8

OpenStudy (crisulcampo):

the part that says 10 chosen lots is wrong... My bad, it should say 10 chosen

OpenStudy (crisulcampo):

@ganeshie8 @agent0smith

OpenStudy (agent0smith):

Pretty confusingly written, but it sounds like a binomial problem. http://fsweb.bainbridge.edu/dbyrd/statistics/binomial_formula.gif N=10 p= 0.05 ("success" is getting an obsolete one) q= 0.95 "lot is always refused if it's found 3 or more obsolete" So it's accepted if there's less than or equal to 2 obsolete ones. You need: \[P(X \le 2) = P(X=0) +P(X=1) +P(X=2) \]

OpenStudy (crisulcampo):

mmmm

OpenStudy (crisulcampo):

I should just replace the values given in link formula, shouldn't I?

OpenStudy (crisulcampo):

I don't know how to apply that formula, sorry bro

OpenStudy (agent0smith):

Yes, i'll show you how to do P(X=0), you can do P(X=1) and P(X=2) the same way. Then just add them all together. \[\large P(X=0) = \left(\begin{matrix}10 \\ 0\end{matrix}\right) 0.05^0 *0.95^{10}\]I assume you know how to do the combinations part; calculators have a button for it, too, or in the math menus.